# A CAV-based perimeter-free regional traffic control strategy utilizing existing parking infrastructure

### v2

@@ 2.2 Maximum stability property
Network throughput is an indicator widely used to assess the performance of traffic control approaches. A control algorithm has the maximum stability property, which is a favorable property investigated extensively for adaptive traffic signal control methods (Varaiya, 2013; Smith et al., 2019), if it maximizes the network throughput, or equivalently, if it can accommodate all feasible demand patterns. This section proves that if the maximum stability property holds for the base control algorithm, it is preserved by the proposed holding strategy.
Definition 1.
@@
Note that, since $x_{(i,j,k)}(t)>0$, Eq. (9) implies that when the network is stabilized, the number of vehicles in the network is always upper bounded. In addition, under the assumption of point queues, which has been commonly employed as an assumption for the proof of maximum stability property (e.g., Varaiya (2013); Smith et al. (2019)), spillover effect is not considered, and the stabilization is independent of the initial state of the network.
Definition 2.
@@
Theorem 1.
[…]
@@
Although the actual values of $\mathbf{N}^{\text{enter}}(t)$ and $\mathbf{N}^{\text{hold}}(t)$ depends on the traffic condition at time $t$, they are always bounded since the parking capacity is assumed to be limited, indicated by Eq. (6). Since control policy $\pazocal{P}$ has the maximum stability property, according to Definition 1, for any $\mathbf{d}\in\mathbf{D}_{f}$, the unconditional expectation of present vehicles from Eq. (11) is bounded by a constant $M^{\prime}$. Additionally, because the total number of held vehicles is always upper bounded by $\sum_{i,j,k}K^{p}_{i,j,k}$, the expectation of total number of vehicles at time $t+1$ is bounded by $M=M^{\prime}+\sum_{i,j,k}K^{p}_{i,j,k}$.
[…]
@@ 3 Numerical Experiment Design
The performance of the proposed control strategy is tested in microscopic traffic simulation using the AIMSUN software. Although the proposed holding strategy is independent of the base traffic signal control strategy, to maintain maximum stability in our tests, we consider the case where all intersections are controlled by the queue-based Max Pressure algorithm proposed by Varaiya (2013), Q-MP, which has the maximum stability property. The holding strategy proposed in the previous section is then applied on Q-MP, called Q-MPH, to show the superiority of the proposed approach on the traffic operational efficiency at a network-level. Both average vehicle delay and MFD — represented by the relationship between vehicle exit rate and the average density in the network — are used to quantify the effectiveness. Since the proposed algorithm introduces a novel paradigm for network-wide traffic management, perimeter control, which has been widely studied for tackling regional congestion, is employed as the baseline method. Specifically, Q-MPH is compared to two perimeter control algorithms: Bang-Bang control and N-MP. The performance of the proposed strategy is investigated in both a pure CAV environment and a mixed CAV-HDV traffic environment. The rest of this section describes the simulation settings and the two benchmark perimeter control algorithms.
@@ 3.1 Simulation settings
We consider here a 10x10 square grid network of two-way streets; see Figure 2. The grid network layout has been widely utilized in traffic flow management for urban transportation systems (Chu et al., 2016; Tan et al., 2019; Chu et al., 2016). The length of all links is set to 200 m. Since Q-MP requires queue lengths for all movements as the input to determine signal timings, we assume that all links have three dedicated lanes, one for each of the left-turn, through, and right-turn movements. The free flow travel speed is 50 km/h, and the saturation for each link is 1,800 vphpl.
@@
    Concentrated demand pattern: In this scenario, travel demand is still generated uniformly throughout the entire network, but all vehicles have destinations within a designated central area. This pattern differs from the uniform demand pattern by removing all trips heading to the outer region. Since perimeter control is designed to mitigate congestion caused by regional high demand, this pattern effectively represents such a scenario and has been widely used in perimeter control studies (Ni and Cassidy, 2019; Liu and Gayah, 2024). Additionally, simulation results indicate that Bang-Bang control performs worse than pure Q-MP control under the uniform demand pattern. Therefore, this pattern is only utilized in Section 4.2.2 for a fair comparison between Q-MPH and the perimeter control methods.
@@ 3.2 Base control: Q-MP
Thanks to its ease of implementation and strong operational performance, MP has become a very popular decentralized traffic signal control algorithm (Wongpiromsarn et al., 2012; Varaiya, 2013; Kouvelas et al., 2014; Xiao et al., 2014; Wu et al., 2017; Li and Jabari, 2019; Dixit et al., 2020; Wang et al., 2022; Mercader et al., 2020; Noaeen et al., 2021; Liu and Gayah, 2022, 2023; Smith and Mounce, 2024; Liu et al., 2024; Ahmed et al., 2024) in the past decade. A comprehensive review of MP algorithms can be found in (Levin, 2023).
The Q-MP (queue-based MP) algorithm proposed by Varaiya (2013) is utilized as the base control for all intersections in the network. To make it easier to follow the proposed method, Q-MP is briefly explained. In Q-MP, the temporal dimension is divided into intervals with fixed duration for signal timing update. At the beginning of each interval, the weight of all movements is calculated as follows:
@@ 3.3 Benchmark perimeter control
Traditional perimeter control algorithms require the identification of a perimeter to limit inflows traversing the perimeter to the congested region. An example of perimeter is shown in Figure 1. Once the perimeter is identified, a general way to limit inflows is to reduce the green time ratio for phases sending vehicles into the congested region at perimeter intersections, i.e., intersections on the perimeter. Two perimeter control approaches, Bang-Bang control (Daganzo, 2007) and N-MP (Liu and Gayah, 2024), are selected as the benchmark perimeter control methods. Both approaches will be explained in more details in the rest of this section. Under the uniform demand pattern defined in Section 3.1, Bang-Bang control results in a worse performance than the baseline pure Q-MP control strategy. Therefore, to ensure a fair comparison, the concentrated demand pattern, which has been utilized to demonstrate the performance of perimeter control methods (Ni and Cassidy, 2019), is employed. Note unlike Bang-Bang control, the performance of the proposed holding control strategy does not rely on this specific demand pattern. Therefore, this modification is only applied in Section 4.2.2, where to ensure a fair comparison between Q-MPH and the two tested perimeter control approaches.
@@ 3.3.1 Bang-Bang control
Bang-Bang control (Daganzo, 2007) is a simple but effective perimeter control strategy, which maximizes the operational efficiency for abstract systems. For this benchmark control strategy, all intersections except for perimeter intersections are still controlled by Q-MP, and Bang-Bang control is incorporated into the Q-MP structure at perimeter intersections only using the following method. At each signal update instant, if the density inside the perimeter, $\rho^{p}(t)$, exceeds the critical density, $\rho^{p}_{cr}$, all inbound lanes at the perimeter intersections will be blocked by adding red time meters, and the signal at perimeter intersections is updated based on Q-MP without considering those inbound lanes. For example, when $\rho^{p}(t)>\rho^{p}_{cr}$, all middle lanes of the northbound approaches at the south-side perimeter intersections (marked by the star symbol in Figure 1) will be blocked and ignored by the weight calculation in Q-MP. If $\rho^{p}(t)<\rho^{p}_{cr}$, perimeter intersections are controlled by the original Q-MP.
@@ 3.3.2 N-MP
Liu and Gayah (2024) recently proposed an algorithm, called N-MP, to incorporate regional perimeter control mechanism directly into MP control structure, while maintaining its decentralized control structure and the maximum stability property. The major difference between N-MP and traditional MP algorithms is that N-MP considers both local and regional traffic states for the weight calculation for the inbound movements at perimeter intersections. When congestion occurs, the weight for the inbound movements decreases with the increase in the density inside the perimeter, $\rho^{p}$, and the increase in the upstream queue length of the local movement. Specifically, the weight for an inbound movement in N-MP is expressed as,
@@ 4 Simulation Results
The performance of the Q-MPH is evaluated in both fully CAV environments – where all vehicles can be requested to hold (Sections 4.1-4.2) – to assess the benefits of the proposed strategy under a complete CAV deployment, and in mixed HDV-CAV environments (Section 4.3) – where only a subset of vehicles can be held – to explore the achievable benefits during the transition era.
@@ 4.1 Results with unrestricted parking space
This section demonstrates the performance of the proposed strategy assuming that parking spaces are available next to all links, as shown in Figure 1, with enough capacity so that CAVs can be held when and where needed. Note that CAVs will be held at the centroid from the same street on which they are currently traveling.
@@ 4.1.2 Impact of holding parameters
Previous results do not impose the threshold on the holding time  [^1] The holding time is set to 5 mins for one-time holding. However, since a CAV can be held multiple times, the total holding time is not limited., which can lead to potential equity issues. To address this issue, we assume that all CAVs can be held at most once. The values of {2, 4, 6, 8, 10} mins are tested for the impact of holding time, $\tau$. The results, shown in Figure 7, reveal that for the tested range, a longer holding time leads to a higher operational efficiency. However, when $\tau\geq 8$ mins, the marginal improvement is less significant than $\tau<8$ mins. Although $\tau=10$ mins generates the best control performance, in practice, the selection for the value of $\tau$ should be determined based on the policy makers’ evaluation on both the improvement for the system’s efficiency and the sacrifice of individual’s time from holding. To avoid holding vehicles for too long, $\tau=6$ mins is used, and all CAVs can be held at most once in the rest of this paper.
@@ 4.1.3 Vehicle travel delay
Figure 8 shows the total delay reduction for both held and non-held vehicles associated with different travel distance. Note that even though the threshold for remaining travel distance is $\phi=8$ links, meaning that vehicles with travel distances less than 8 links will never be held, there are held vehicles that exist with total travel distances less than 8 blocks in Figure 8. This occurs because these travel distances indicated on the horizontal axis of Figure 8 represent the travel distance under the Q-MP (i.e., without holding); due to randomness and the dynamic routing used in the simulation, these vehicles might sometime take a route longer than $\phi$ blocks when holding is applied. As a result, they can be held and their travel delay will be increased, indicating by the negative delay reductions in Figure 8. However, the percentage of these vehicles is very small, shown in Figure 9.
In line with our expectation, the travel delay for all non-held vehicle groups, shown by the dashed red line in Figure 8, is reduced. More interestingly, we notice that although the travel delay of held-vehicles with a travel distance between 8-12 links increases, the travel delay for the held-vehicle groups with a longer travel distance actually decreases. This indicates that the holding strategy does not only save time for non-held vehicles, but it can also be beneficial for those CAVs that are held, leading to a reduction for total travel delay for all vehicle groups. Assume a traffic system in which all travelers’ ODs are fixed, but the travel time is flexible from day to day, the delay reduction for all vehicle groups implies that everyone can benefits from the proposed strategy in the long run, which serves as a favorable feature of the strategy.
@@
Figure 10 shows the average delay reduction for both held and non-held vehicles. Interestingly, it indicates that for both held and non-held vehicles, the proposed control strategy brings greater advantages (lower disadvantages for held vehicles with delay increase) for vehicles with longer travel distances. It is reasonable to assume that, on average, the travel time reduction for a unit travel distance generated from the proposed strategy is identical for all vehicles. As a result, the total delay reduction increases with travel distance.
@@
Figure 9: Percentage of vehicle groups.
Figure 10: Average vehicle delay reduction.
@@ 4.2 Results with restricted parking space
Figure 12 shows the results associated with the five parking location patterns in Figure 11. Overall, $i=7$ generates the highest operational efficiency among the tested values. When $i=9$, although more parking locations are offered, the number of trips passing those locations is smaller than that when $i=7$. Consequently, fewer vehicles will be held, and the enhancement of the efficiency is less significant. On the other hand, a smaller value for $i$ can also lead to a reduction in the control performance for two reasons: 1) the number of parking locations is reduced; 2) the number of vehicles passing those locations that have a remaining travel distance longer than the threshold value is also reduced. For example, if we assume all vehicles are always using distance-based shortest path to their destination, when $i=1$, the longest remaining travel distance for vehicles passing any of the four centroids is 10 links. However, this maximum value for a centroid when $i=7$ is 16 links, which leads to a higher number of vehicles that can be potentially held.
@@
4.2.1 Results with parking capacity limits

### v3

@@ 2.3 Maximum stability property
Network throughput is an indicator widely used to assess the performance of traffic control approaches. A control algorithm has the maximum stability property, which is a favorable property investigated extensively for adaptive traffic signal control methods [46, 40], if it maximizes the network throughput, or equivalently, if it can accommodate all feasible demand patterns. This section proves that if the maximum stability property holds for the base control algorithm, it is preserved by the proposed holding strategy.
Definition 2.1.
@@
Note that, since $x_{(i,j,k)}(t)>0$, Eq. (12) implies that when the network is stabilized, the number of vehicles in the network is always upper bounded. In addition, under the assumption of point queues, which has been commonly employed as an assumption for the proof of maximum stability property (e.g., [46, 40]), spillover effect is not considered, and the stabilization is independent of the initial state of the network.
Definition 2.2.
@@
Theorem 2.4.
If policy $\pazocal{P}$ has the maximum stability and the maximum number of holding vehicles $x_{\text{max}}^{\text{CAV,H}}$ in the network is bounded, the incorporation of the proposed holding strategy into $\pazocal{P}$, denoted by $\pazocal{P}^{\textsuperscript{H}}$, has the maximum stability property.
@@
Although the actual values of $\mathbf{N}^{\text{enter}}(t)$ and $\mathbf{N}^{\text{hold}}(t)$ depends on the traffic condition at time $t$, they are always bounded since the parking capacity is assumed to be limited, indicated by Eq. (6). Since control policy $\pazocal{P}$ has the maximum stability property, according to Definition 2.1, for any $\mathbf{d}\in\mathbf{D}_{f}$, the unconditional expectation of present vehicles from Eq. (14) is bounded by a constant $M^{\prime}$. Additionally, because the total number of held vehicles is always upper bounded by $\sum_{i,j,k}K^{p}_{i,j,k}$, the expectation of total number of vehicles at time $t+1$ is bounded by $M=M^{\prime}+\sum_{i,j,k}K^{p}_{i,j,k}$.
Above all, if $\mathbf{1}^{\top}\mathbf{x}(1)$ is bounded, the evolution of number of vehicles in the network resulting from the control policy $\pazocal{P}^{\textsuperscript{H}}$ satisfies (12) for any $\mathbf{d}\in\mathbf{D}_{f}$. As a result, policy $\pazocal{P}^{\textsuperscript{H}}$ has the maximum stability property.
@@ 3 Numerical Experiment Design
The performance of the proposed control strategy is tested in microscopic traffic simulation using the AIMSUN software. Although the proposed holding strategy is independent of the base traffic signal control strategy, to maintain maximum stability in our tests, we consider the case where all intersections are controlled by the queue-based Max Pressure algorithm proposed by [46], Q-MP, which has the maximum stability property. The holding strategy proposed in the previous section is then applied on Q-MP, called Q-MPH, to show the superiority of the proposed approach on the traffic operational efficiency at a network-level. Both average vehicle delay and MFD — represented by the relationship between vehicle exit rate and the average density in the network — are used to quantify the effectiveness. Since the proposed algorithm introduces a novel paradigm for network-wide traffic management, perimeter control, which has been widely studied for tackling regional congestion, is employed as the baseline method. Specifically, Q-MPH is compared to two perimeter control algorithms: Bang-Bang control and N-MP. The performance of the proposed strategy is investigated in both a pure CAV environment and a mixed CAV-HDV traffic environment. The rest of this section describes the simulation settings and the two benchmark perimeter control algorithms.
@@ 3.1 Simulation settings
We consider here a 10x10 square grid network of two-way streets; see Figure 1. The grid network layout has been widely utilized in traffic flow management for urban transportation systems [6, 43, 6]. The length of all links is set to 200 m. Since Q-MP requires queue lengths for all movements as the input to determine signal timings, we assume that all links have three dedicated lanes, one for each of the left-turn, through, and right-turn movements. The free flow travel speed is 50 km/h, and the saturation for each link is 1,800 vphpl.
Figure 1: Network setup.
@@
    Concentrated demand pattern: In this scenario, travel demand is still generated uniformly throughout the entire network, but all vehicles have destinations within a designated central area. This pattern differs from the uniform demand pattern by removing all trips heading to the outer region. Since perimeter control is designed to mitigate congestion caused by regional high demand, this pattern effectively represents such a scenario and has been widely used in perimeter control studies [32, 28]. Additionally, simulation results indicate that Bang-Bang control performs worse than pure Q-MP control under the uniform demand pattern. Therefore, this pattern is only utilized in Section 4.2.3 for a fair comparison between Q-MPH and the perimeter control methods.
@@ 3.2 Base control: Q-MP
Thanks to its ease of implementation and strong operational performance, MP has become a very popular decentralized traffic signal control algorithm [51, 46, 17, 54, 52, 20, 8, 49, 30, 33, 26, 27, 41, 25, 1] in the past decade. A comprehensive review of MP algorithms can be found in [19].
The Q-MP (queue-based MP) algorithm proposed by [46] is utilized as the base control for all intersections in the network. To make it easier to follow the proposed method, Q-MP is briefly explained. In Q-MP, the temporal dimension is divided into intervals with fixed duration for signal timing update. At the beginning of each interval, the weight of all movements is calculated as follows:
@@ 3.3 Benchmark perimeter control
Traditional perimeter control algorithms require the identification of a perimeter to limit inflows traversing the perimeter to the congested region. An example of perimeter is shown in Figure 1. Once the perimeter is identified, a general way to limit inflows is to reduce the green time ratio for phases sending vehicles into the congested region at perimeter intersections, i.e., intersections on the perimeter. Two perimeter control approaches, Bang-Bang control [7] and N-MP [28], are selected as the benchmark perimeter control methods. Both approaches will be explained in more details in the rest of this section. Note that [28] has showed that the feedback-based perimeter control [15] generated worse performance than Bang-Bang control under similar settings. Therefore, the former is excluded from this paper.
Under the uniform demand pattern defined in Section 3.1, Bang-Bang control results in a worse performance than the baseline pure Q-MP control strategy. Therefore, to ensure a fair comparison, the concentrated demand pattern, which has been utilized to demonstrate the performance of perimeter control methods [32], is employed. Note unlike Bang-Bang control, the performance of the proposed holding control strategy does not rely on this specific demand pattern. Therefore, this modification is only applied in Section 4.2.3, where to ensure a fair comparison between Q-MPH and the two tested perimeter control approaches.
@@ 3.3.1 Bang-Bang control
Bang-Bang control [7] is a simple but effective perimeter control strategy, which maximizes the operational efficiency for abstract systems. For this benchmark control strategy, all intersections except for perimeter intersections are still controlled by Q-MP, and Bang-Bang control is incorporated into the Q-MP structure at perimeter intersections only using the following method. At each signal update instant, if the density inside the perimeter, $\rho^{p}(t)$, exceeds the critical density, $\rho^{p}_{cr}$, all inbound lanes at the perimeter intersections will be blocked by adding red time meters, and the signal at perimeter intersections is updated based on Q-MP without considering those inbound lanes. For example, when $\rho^{p}(t)>\rho^{p}_{cr}$, all middle lanes of the northbound approaches at the south-side perimeter intersections (marked by the star symbol in Figure 1) will be blocked and ignored by the weight calculation in Q-MP. If $\rho^{p}(t)<\rho^{p}_{cr}$, perimeter intersections are controlled by the original Q-MP.
@@ 3.3.2 N-MP
[28] recently proposed an algorithm, called N-MP, to incorporate regional perimeter control mechanism directly into MP control structure, while maintaining its decentralized control structure and the maximum stability property. The major difference between N-MP and traditional MP algorithms is that N-MP considers both local and regional traffic states for the weight calculation for the inbound movements at perimeter intersections. When congestion occurs, the weight for the inbound movements decreases with the increase in the density inside the perimeter, $\rho^{p}$, and the increase in the upstream queue length of the local movement. Specifically, the weight for an inbound movement in N-MP is expressed as,
@@ 4 Simulation Results
The performance of the Q-MPH is evaluated in both fully CAV environments – where all vehicles can be requested to hold (Sections 4.1-4.2) – to assess the benefits of the proposed strategy under a complete CAV deployment, and in mixed HDV-CAV environments (Section 4.3) – where only a subset of vehicles can be held – to explore the achievable benefits during the transition era. In addition, to assess the feasibility of the proposed methods, various configurations of parking facilities—including number, distribution, and capacity—are tested in the simulation. Moreover, the effectiveness of holding long-distance trips is further evaluated under a more realistic setting characterized by irregular spatial parking distributions and limited parking capacity. This setup allows us to control for differences in the number of vehicles removed from the network, thereby isolating the effect of the holding rule itself. The network exit rate, a commonly used metric for evaluating regional traffic management performance, is used as the primary performance index. Furthermore, we examine the impact of the proposed method on vehicle delays for both held and non-held vehicles. The scenarios, in terms of the spatial distribution and capacity of parking facilities, in the rest of this section are summarized below:
  • Section 4.1: parking available to all links with infinite parking capacity
  • Section 4.2: restricted parking space
    • Section 4.2.1: perimeter-like parking distribution with infinite parking capacity
    • Section 4.2.2: perimeter-like parking distribution with limited parking capacity
    • Section 4.2.3: perimeter-like parking distribution with infinite parking capacity
    • Section 4.2.4: random parking distribution with infinite parking capacity
    • Section 4.2.5: random parking distribution with limited parking capacity
  • Section 4.3: random parking distribution with infinite parking capacity under varying CAV penetration rates
@@ 4.1 Results with unrestricted parking space
This section demonstrates the performance of the proposed strategy assuming that parking spaces are available next to all links, as shown in Figure 1, with enough capacity so that CAVs can be held when and where needed. This scenario is employed to demonstrate the efficiency of the proposed strategy under an idealized setting. Note that CAVs will be held at the centroid from the same street on which they are currently traveling.
@@ 4.1.2 Impact of holding parameters
The previous results do not impose a threshold on the holding time  [^2] The holding time is set to 5 mins for one-time holding. However, since a CAV can be held multiple times, the total holding time is not limited., which can lead to potential equity issues. To address this issue, we assume that all CAVs can be held at most once and vary the amount of time a vehicle is held, $\tau$. The results, shown in Figure 7, reveal that for the tested range, a longer holding time leads to a higher operational efficiency. However, when $\tau\geq 8$ mins, the marginal improvement is less significant than $\tau<8$ mins. Although $\tau=10$ mins generates the best control performance, in practice, the selection for the value of $\tau$ should be determined based on the policy makers’ evaluation on both the improvement for the system’s efficiency and the sacrifice of individual’s time from holding. To avoid holding vehicles for too long, $\tau=6$ mins is used, and all CAVs can be held at most once in the rest of this paper.
@@ 4.1.3 Vehicle travel delay
Figure 8 shows the total delay reduction for both held and non-held vehicles associated with different travel distance. Note that even though the threshold for remaining travel distance is $\phi=8$ links, meaning that vehicles with travel distances less than 8 links will never be held, there still exists held vehicleswith total travel distances less than 8 blocks in Figure 8. This occurs because these travel distances indicated on the horizontal axis of Figure 8 represent the travel distance under the Q-MP (i.e., without holding); due to randomness and the dynamic routing used in the simulation, these vehicles might sometime take a route longer than $\phi$ blocks when holding is applied. As a result, they can be held and their travel delay will be increased, indicating by the negative delay reductions in Figure 8. However, the percentage of these vehicles is very small as to be negligible; see Figure 9.
@@
Figure 9: Percentage of vehicle groups.
In line with our expectation, travel delays are reduced for all vehicles that are not held; see the dashed red line in Figure 8. Perhaps more importantly, however, we notice that although the travel delay of held-vehicles with a travel distance between 8-12 links increases, the travel delay for the held-vehicle groups with a longer travel distance actually decreases. This indicates that the holding strategy not only saves time for non-held vehicles, but it can also potentially benefit those CAVs that are held, leading to a reduction for total travel delay for all vehicle groups. This occurs because traffic conditions under the proposed strategy are improved compared to those that would arise without it, and longer trips tend to receive greater benefits from the reduced congestion. The observed delay reductions across these vehicle groups suggest that users in these groups can benefit from the long run, which represents a desirable property. This is an important feature of the proposed approach, as human drivers may also be more willing to participate if they recognize that temporary waiting can ultimately reduce their own travel times, thereby enhancing the practical implementation of the strategy.
Figure 10 shows the average delay reduction per trip for both held and non-held vehicles. Interestingly, it indicates that for both held and non-held vehicles, the proposed control strategy brings greater advantages (or lower disadvantages for held vehicles that experience a delay increase) for vehicles with longer travel distances. It is reasonable to assume that, on average, the travel time reduction for a unit travel distance generated from the proposed strategy is identical for all vehicles. As a result, the total delay reduction increases with travel distance.
Figure 10: Average vehicle delay reduction.
@@
4.1.4 Impact of delay from entering/exiting parking facilities
Although parking facilities are associated with links such that CAVs do not need to cruise or reroute to search for parking, under congested conditions, additional time may still be required for vehicles to fully exit the network after initiating the parking maneuver. Therefore, we conduct an additional sensitivity analysis for the scenario with $\phi=8$ links and $\tau=6$ min, in which the enter/exit delay is modeled as a normally distributed random variable with mean $\mu_{d}$ ranging from 10 s to 60 s and a standard deviation of 5 s. The effective holding time is adjusted by subtracting this delay from the predefined holding duration $\tau=6$ min. If a sampled delay is negative, it is set to zero; similarly, if the adjusted holding time becomes negative, it is truncated to zero. The impact of this delay on control performance is illustrated in Figure 11, which shows that the strategy remains robust across the tested scenarios. As the enter/exit delay increases, the number of exiting vehicles and the total delay reduction decrease slightly, reflecting the reduced effective holding time. However, this impact appears to be negiglible on the overall operating performance.
(a) Comparison of exit vehicles
(b) Vehicle delay
Figure 11: Influence of $\mu_{d}$ under $\phi=8$ links, $\tau=6$ mins.
@@ 4.2 Results with restricted parking space
4.2.1 Perimeter-like parking distribution with infinite parking capacity
@@
Figure 13 shows the results associated with the five parking location patterns in Figure 12. Overall, $i=7$ generates the highest operational efficiency among the tested values. When $i=9$, although more parking locations are offered, the number of trips passing those locations is smaller than that when $i=7$. Consequently, fewer vehicles may be held, and the enhancement of the efficiency is less significant. On the other hand, a smaller value for $i$ can also lead to a reduction in the control performance for two reasons: 1) the number of parking locations is reduced; 2) the number of vehicles passing those locations that have a remaining travel distance longer than the threshold value is also reduced. For example, if we assume all vehicles are always using distance-based shortest path to their destination, when $i=1$, the longest remaining travel distance for vehicles passing any of the four centroids is 10 links. However, this maximum value for a centroid when $i=7$ is 16 links, which leads to a higher number of vehicles that can be potentially held.
@@
4.2.2 Perimeter-like parking distribution with parking capacity limits